Calculate the equilibrium concentrations of N2O4 and NO2 at 25 ∘C if the initial concentrations are [N2O4]= 0.0130 M and [NO2]= 0.0300 M. The equilibrium constant Kc for the reaction N2O4(g)⇌2NO2(g) is 4.64×10−3 at 25 ∘C. I've worked through this problem a few times and still don't get the right answer. Can anyone show me the work??
Answer #1

N28469) 2248219) kc = 4.64x10-3 initial concentration of CM204] = 0.0130 m [MO2] = 0.0380m 2 Sm using ICE table N20419) Z 2N8ux² +0.12% + 4.64 X10-3x + 9x10-4_6.032x10-5=0 4x² +0.124x + 8.40 X10-9=o x² + 0.031 +2.1x18 =o it is a quadratic equation (